# A Novel Passive Occupational Shoulder Exoskeleton With Adjustable Peak Assistive Torque Angle For Overhead Tasks

### v1

@@ 4.1 ROM
Table 5 reports measured ROM averaged among all participants. No significant differences were observed between conditions. The absolute error between the maximum value (169.71°) of the penultimate set in the simulation of Fig. 5 and the maximum sagittal F/E angle (164.46°) in HIT-POSE condition of the experiment was 5.25°. Fig. 13 shows the averaged hand workspace (normalized by the arm length of each subject) of the maximum sagittal and horizontal F/E angles among all participants in both conditions. In both movements, the coordinate origin is the hand’s position in the neutral posture before movements. The positive x-axis points to the right side of the body, the positive y-axis points forward, and the positive z-axis points upward. There was no significant statistical difference between conditions for both tasks.
Figure 13: Averaged hand workspace (normalized by the arm length of each subject) of the maximum sagittal and horizontal F/E angles among all participants in both conditions. (a) Hand workspace of sagittal F/E, (b) Hand workspace of horizontal F/E. The error bar represents the mean ± standard deviation.
@@ 4.2 Muscular Activations
For all monitored muscles, statistically significant interactions were found in both conditions, as shown in Table 6. Fig. 14 demonstrates the results of muscular activations from pairwise comparisons. In HIT-POSE Match conditions for both tasks, significantly decreased muscular activations in all monitored muscles were found compared to NO HIT-POSE condition. For screwing at a low height, the muscular activations of MD, PD, BB, PM, and LD muscles in HIT-POSE Match condition were significantly reduced compared to that in HIT-POSE Mismatch condition; in addition, the muscular activation of the BB muscle in HIT-POSE Mismatch condition was also significantly reduced compared to that in NO HIT-POSE condition. For screwing at a high height, AD, MD, PD, and PM muscles in HIT-POSE Match condition exhibited significantly decreased EMG activations compared to that in HIT-POSE Mismatch condition; moreover, BB and TR muscles in HIT-POSE Mismatch condition exhibited significantly decreased EMG activations compared to that in NO HIT-POSE condition.
[…]
Table 6: F AND P-VALUES OF THE ANOVAS ON THE MUSCULAR ACTIVATIONS DURING THE TWO TASKS.
Muscles Screwing at a low height Screwing at a high height
F-value P-value F-value P-value
AD 11.65 0.0006 13.03 0.0003
MD 16.84 0.0001 12.33 0.0004
PD 18.75 0.0000 19.91 0.0000
BB 13.75 0.0002 13.25 0.0003
PM 11.23 0.0007 25.60 0.0000
TR 10.08 0.0012 11.85 0.0005
LD 10.69 0.0009 8.10 0.0031
ES 12.00 0.0005 11.61 0.0006
@@ 4.3 Subjective Feedback
For RPE scores, statistically significant interactions were found in both conditions (F = 14.58, p = 0.0002 for low height; F = 10.55, p = 0.0009 for high height). Fig. 15 shows the RPE scores averaged across participants from pairwise comparisons. In HIT-POSE Match conditions for both tasks, RPE scores showed a significant difference relative to the other two conditions. For physical demand, effort, and frustration aspects of NASA-TLX, statistically significant interactions were found in both conditions (Physical Demand: F = 5.52, p = 0.0135 for low height; F = 5.08, p = 0.0179 for high height. Effort: F = 7.76, p = 0.0037 for low height; F = 6.85, p = 0.0061 for high height. Frustration: F = 6.34, p = 0.0082 for low height; F = 4.08, p = 0.0345 for high height). For both tasks in HIT-POSE Match conditions, significant differences were found in physical demand, effort, and frustration compared to NO HIT-POSE conditions (Fig. 16). In addition, the average SUS score was 79.7 ± 5.9.
@@ 5.1 Kinematics Compatibility
HIT-POSE is equipped with an ergonomic shoulder structure to ensure sufficient ROM for the user. This implies favorable kinematics compatibility, which has a profound impact on the exoskeleton’s usability and user acceptance. Typically, some passive shoulder exoskeletons integrate a multi-DoF chain or simplified two-link hinge to increase the ROM, such as the SPM [23] and HUST-EV [46]. However, they may collide with nearby objects easily due to the large volume and protrusions. In addition, the design parameters were fixed, thus lacking reconfigurability. This reduces the adaptability and increases the manufacturing costs of the exoskeleton.
To ensure sufficient ROM while compactness, we extracted three parameters ($\phi$, $d_{v}$, $d_{b}$) to design the shoulder structure, and two of them ($d_{v}$ and $d_{b}$) an be adjusted without reconstructing the exoskeleton. A suitable set of parameters ($\phi$=15°, $d_{v}$=80mm, $d_{b}$=10mm) was selected through kinematic modeling and ROM simulations. Constrained but sufficient ROM was guaranteed in both sagittal and horizontal F/E for performing overhead work, as shown in Table 5, while the structure with a relatively small volume (22cm × 48cm × 15cm) is designed to reduce the possibility of collision. Since volume information for other devices was not found, no comparison was made. However, our exoskeleton volume is visually more compact than its counterparts. The discrepancies between the ROM in simulations and experiments may be due to the omission of friction and contact forces in the simulations.
We had a trade-off for the shoulder movements, limiting the abduction/adduction movements of the shoulder, which are not the focus of overhead work. Through the coordination of sagittal and horizontal F/E joints, the exoskeleton can achieve abduction/adduction movements in the anterior space of the body, but there are limitations on movements in the posterior area of the body, such as extending the elbow diagonally backward and outward. This method meets the design requirements of the overhead work, and similar degree of freedom configuration strategies can be found on the ShoulderX [47] and Airframe [22]. As shown in Fig. 13, wearing the exoskeleton reduced hand workspace, likely due to a corresponding reduction in shoulder ROM.
@@ 5.2 Comparison of Adaptability with Existing Exoskeletons
Various overhead tasks and individuals usually have different PATAs and thus different optimal torque profiles [25], [26]. Consequently, the task and individual adaptability of the exoskeleton are crucial, which essentially involves adjustable PATA and peak torque. Existing exoskeletons have made considerable efforts in adjusting the assistive levels with different peak torque. Several exoskeletons employ a manual knob or an electrical component to adjust the assistive level, such as the Shoulder X [21] and H-PULSE [9, 13], other devices require to replace the spring or other parts affecting the spring length. However, existing exoskeletons can only adjust the amplitude of peak torque, thus just adapting to a specific task. Employing torque profiles with a fixed PATA to cope with all different tasks and individuals is not an energy-efficient way and limits the adaptability of the torque generator [26], [31]. For such exoskeletons, if we need to adjust the PATA, we can only change it by reconstructing the exoskeleton.
To enhance the task adaptability and personalized customization capability of the exoskeleton, we proposed a torque generator with an adjustable PATA. We utilized a set of parameters ($\alpha$, $\beta$, $r_{1}$, $r_{2}$, $r_{3}$, $L_{i}$) to guide the design of the torque generator, and investigated their impacts on torque profiles via geometric analysis and simulations. A subset of parameters ($\alpha$, $\beta$, $L_{i}$) can be adjusted to reconfigure the exoskeleton, each controlling peak torque, PATA, and safety limit angle, respectively. Notably, $\beta$ can be adjusted within a range of -30º to 30º, resulting in a range of PATA from 90º to 150º, which meets the requirements of various overhead tasks and individuals. The experimental results demonstrated that monitored muscular activations in HIT-POSE Match condition decreased the most compared to the NO HIT-POSE condition, thus validating the effectiveness of PATA. However, the exoskeleton provided an assistive torque of around 2 N·m when lowering the arm, which may cause discomfort to users. This is inevitable in our current design and also a limitation of the current architecture. Although this is not reflected in users’ subjective evaluations, it may be attributed to the small sample size and individual sensitivity.
@@ 5.3 Effects of the Exoskeleton on Muscular Activations
For both tasks in HIT-POSE Match conditions, significantly reduced muscular activations were observed in agonist muscles (AD, MD, BB, PM, TR) and antagonist muscles (PD, LD) compared to NO HIT-POSE conditions. The exoskeleton provided support to the human shoulder, allowing agonist muscles to generate less force to achieve the same action, while also decreasing co-contraction of the antagonist muscles [13]. Similar results can be found in the ShoulderX [21], PAEXO [48], MATE [32], and H-VEX [26]. This phenomenon can be explained by the fact that there is a match between the PATA and the target task angle, resulting in increased assistive torque, with reductions consistent with the level of assistance [13, 49]. This also demonstrates the necessity of PATA for effective assistance. As reported by previous studies [50, 51], the reduced muscular activation implies lower force, resulting in decreased compressive force on the shoulder, and thus potentially reducing biomechanical stress on the shoulder girdle. Such positive effects suggest that long-term use of the exoskeleton has the potential to reduce the incidence of WMSDs, as reported by previous studies [17], [24].
Additionally, for screwing tasks in HIT-POSE Match conditions, the primary muscles (AD, MD, PD, BB, PM, and LD) significantly reduced their activations compared to HIT-POSE Mismatch conditions. The phenomenon is likely since a larger assistance was provided in HIT-POSE Match conditions, which provides evidence of our rational design. Moreover, AD muscle in screwing at a low height did not exhibit a significant decrease between the two conditions. A previous study has shown that the activation reduction of AD muscle in high tasks is greater than that in low tasks [11], which may be a possible reason. Furthermore, the muscular activations of BB and TR muscles in screwing at a high height did not significantly change between the two conditions. In screwing at a high height, we observed that most participants involved more arm extension (except for lowering the arm), compared to screwing at a low height. The mechanical hysteresis of the spring contributes to larger assistance involving arm extension, which may preferentially transmit to BB muscle.
Notably, compared to NO HIT-POSE conditions, reductions in ES muscle were also expected for both tasks in HIT-POSE Match conditions. This behavior is consistent with the existing study [52]. This may be because the exoskeleton transmits forces to the waist through the device’s back bar and waist belt. This result could indicate that HIT-POSE doesn’t increase the risk of back WMSDs. An unexpected phenomenon is that the muscular activations in Fig. 14 were quite high. This could be attributed to our different normalization method, which does not affect the trend of EMG changes.
@@ 5.4 Effects of The Exoskeleton on Subjective Feedback
For both tasks in HIT-POSE Match conditions, the RPE displayed statistical significance from the other two conditions. Compared to NO HIT-POSE conditions, scores of the physical demand and effort items exhibited a significant reduction for both tasks in HIT-POSE Match conditions. Participants perceived the greatest assistance and thus exerted the lowest physical effort. This outcome is in line with the existing study [13]. In addition, for both tasks in HIT-POSE Match conditions, scores of the frustration item also exhibited a significant reduction with NO HIT-POSE conditions, which is consistent with the previous study [24]. Evidence may be available in a SUS question (9. I felt very confident using the system.) with a high average score (4.56 ± 0.53). We speculated that mental confidence bridges the frustration from the task, which may be related to a placebo effect. Meanwhile, the SUS score (79.7 ± 5.9) expressed good usability of the exoskeleton according to the empirical evaluation [53].
@@ 6 CONCLUSION
In the future, we will strive to enhance the kinematic compliance of the exoskeleton to achieve a full ROM for the user. More realistic simulation algorithms of ROM will also be investigated to reduce the deviation between simulation and reality. In addition, a unique cam structure will be designed to further reduce the assistive torque at low elevation angles and improve the user acceptance and usability of the exoskeleton.

### v2

@@ 4.1 ROM
Table 5 reports measured ROM averaged among all participants in NO HIT-POSE and HIT-POSE conditions. No significant differences were observed in maximum sagittal and horizontal F/E angle between conditions ( p = 0.6432 and p = 0.7421, respectively).
@@ 4.2 Muscular Activations
We evaluated the impact of the exoskeleton on muscular activations, by comparing wearing and not wearing the exoskeleton and comparing Match and Mismatch conditions. Herein, “wearing the HIT-POSE” refers to both HIT-POSE Match and HIT-POSE Mismatch conditions. As shown in Fig. 13, when wearing the HIT-POSE to perform both tasks, the AD, MD, PD, BB, TR, and ES muscles exhibited significantly decreased EMG activations compared to not wearing the HIT-POSE. In HIT-POSE Match conditions for both tasks, significantly decreased muscular activations in all monitored muscles were found compared to NO HIT-POSE and HIT-POSE Mismatch conditions. In HIT-POSE Mismatch conditions for both tasks, compared to the NO HIT-POSE condition, there is a significant reduction in muscle activation for all monitored muscles other than PM and LD.
As shown in Table 6, in HIT-POSE Match conditions for both tasks, the absolute percentage reduction was above 15% (highest at 22.9% and 23.8% for AD in screwing at low and high heights, respectively), and the relative percentage reduction was above 30% (highest at 48.3% for AD in screwing at a low height and 49.6% for BB in screwing at a high height). In HIT-POSE Mismatch conditions for both tasks, the absolute percentage reduction was above 5% (highest at 11.9% for AD in screwing at a low height, and 11.3 for BB and TR in screwing at a high height), and the relative percentage reduction was above 9% (highest at 24.3% for AD in screwing at a low height and 27.7% for BB in screwing at a high height). For the same muscle in the same task, the absolute and relative decrease percentages in the HIT-POSE Match condition are approximately twice as much as in the HIT-POSE Mismatch condition.
@@ 4.3 Subjective Feedback
Fig. 14 shows the RPE scores averaged across participants. In HIT-POSE Match conditions for both tasks, RPE scores showed a significant difference relative to the other two conditions. In HIT-POSE Mismatch conditions for both tasks, no significant differences were observed compared to the NO HIT-POSE condition. NASA-TLX questionnaire assesses PW across six aspects. As indicated in Fig. 15, in HIT-POSE Match conditions for both tasks, significant differences were found in physical demand, effort, and frustration, compared to the other two conditions. No significant differences were observed in mental demand, temporal demand, and performance. In addition, the average SUS score was 79.7 ± 5.9.
@@ 5 DISCUSSION
In this study, we propose a novel passive occupational shoulder exoskeleton with adjustable PATA for different overhead tasks. We first designed an ergonomic shoulder structure through kinematic modeling and ROM simulations to ensure a sufficient shoulder ROM. Then we implemented a torque generator with pulleys moved along the guide rail, thus realizing adjustable PATA. Finally, we also investigated parameters’ impact on torque profiles, providing users with valuable recommendations for achieving various torque profiles through the considerable reconfiguration flexibility of the exoskeleton to adapt to diverse tasks and individuals.
@@ 5.1 Kinematics Compatibility
The HIT-POSE is equipped with an ergonomic shoulder structure to ensure sufficient ROM for the user. This implies favorable kinematics compatibility, which has a profound impact on the exoskeleton’s usability and user acceptance. Typically, some passive shoulder exoskeletons integrate a multi-DoF chain or simplified two-link hinge to increase the ROM, such as MATE [37] and PAEXO [38]. However, they may collide with nearby objects easily due to the large volume and protrusions. Recently, a similar study has proposed a shoulder structure to assist overhead tasks in the shipbuilding industry [11]. However, the ROM of the horizontal F/E is limited, and the design parameters are fixed, thus lacking reconfigurability. This reduces the adaptability and increases the manufacturing costs of the exoskeleton.
To ensure sufficient ROM while compactness, we extracted three parameters ($\phi$, $d_{v}$, $d_{b}$) to design the shoulder structure, and two of them ($d_{v}$ and $d_{b}$) an be adjusted without reconstructing the exoskeleton. A suitable set of parameters ($\phi$=15°, $d_{v}$=80mm, $d_{b}$=10mm) was selected through kinematic modeling and ROM simulations.Sufficient ROM was guaranteed in both sagittal and horizontal F/E for performing overhead tasks and natural movements, as shown in Table LABEL:5, while the structure with a relatively small volume is designed to reduce the possibility of collision.
@@ 5.2 Comparison of Adaptability with Existing Exoskeletons
Different overhead tasks usually have different PATAs and thus different optimal torque profiles [27, 22]. Additionally, when handling the same task, individuals with different heights may have different target angles, i.e., different tasks. Consequently, the task and individual adaptability of the exoskeleton are crucial for users, which essentially involves adjustable assistive profiles, including peak torque and PATA. Existing exoskeletons have made considerable efforts in adjusting the assistive levels with different peak torque. Several exoskeletons employ a manual knob or an electrical component to adjust the assistive level, such as the Shoulder X [17] and H-PULSE [9, 12], other devices require to replace the spring or other parts affecting the spring length. However, existing exoskeletons can only adjust the amplitude of peak torque, thus just adapting to a specific task. Employing torque profiles with a fixed PATA to cope with all different tasks and individuals is not an energy-efficient way and limits the adaptability of the torque generator [22], [28]. For such exoskeletons, if we need to adjust the PATA, we can only change it by reconstructing the exoskeleton.
To enhance the task adaptability and personalized customization capability of the exoskeleton, we proposed a torque generator with an adjustable PATA. We utilized a set of parameters ($\alpha$, $\beta$, $r_{1}$, $r_{2}$, $r_{3}$, $L_{i}$) to guide the design of the torque generator, and investigated their impacts on torque profiles via geometric analysis and simulations. A subset of parameters ($\alpha$, $\beta$, $L_{i}$) can be adjusted to reconfigure the exoskeleton, each controlling peak torque, PATA, and safety limit angle, respectively. Notably, $\beta$ can be adjusted within a range of -30º to 30º, resulting in a range of PATA from 90º to 150º, which meets the requirements of various overhead tasks and individuals. The experimental results demonstrated that monitored muscular activations in HIT-POSE Match condition decreased the most compared to the NO HIT-POSE condition, thus validating the effectiveness of PATA.
@@ 5.3 Effects of the Exoskeleton on Muscular Activations
To provide a comprehensive overview of the effects of the HIT-POSE on muscular activations, we monitored and analyzed eight muscles primarily involved in shoulder movement and stability in three conditions (i.e., NO HIT-POSE, HIT-POSE Match, and HIT-POSE Mismatch) during screwing at low and high heights. Overall, our findings indicated that the HIT-POSE can significantly reduce muscular activations of related muscles, thus potentially mitigating the incidence of shoulder WMSDs.
When wearing the exoskeleton in both tasks, significantly reduced muscular activations were observed in agonist muscles (AD, MD, BB, TR) and antagonist muscles (PD). The exoskeleton provided support to the human shoulder, allowing agonist muscles to generate less force to achieve the same action, while also decreasing co-contraction of the antagonist muscles [12]. Similar results can be found in the ShoulderX [39], PAEXO [40], MATE [25], and H-VEX [22]. This observation reflects the rational and effective design of the device. As reported by a previous study [41], the reduced muscular activation implies lower force, resulting in decreased compressive force on the shoulder, and thus potentially reducing biomechanical stress on the shoulder girdle. Such positive effects suggest that long-term use of the exoskeleton has the potential to reduce the incidence of WMSDs, as reported by previous studies [15, 26]. Additionally, for both tasks, PM (agonist muscle) and LD (antagonist muscle) demonstrated significantly reduced muscular activations in the HIT-POSE Match condition, but not in the HIT-POSE Mismatch condition. The difference is likely due to the fact that less assistance was provided in the HIT-POSE Mismatch condition, or individuals adopted different shoulder postures for different tasks as well as conditions. Nevertheless, the more accurate explanations remain to be explored further.
Notably, a reduction in the ES muscle was also expected when wearing the HIT-POSE in both tasks. The result is consistent with the existing study [42]. This may be attributed to the fact that the exoskeleton transmits forces to the waist through the device’s back bar and waist belt. The result could indicate that the HIT-POSE doesn’t increase the risk of back WMSDs. As expected, for both tasks, all muscular activations decreased the most in the HIT-POSE Match condition. This phenomenon can be explained by the fact that matching the PATA with the target task angle results in increased assistive torque, with reductions consistent with the level of assistance [12]. This also demonstrates the necessity of PATA for effective assistance.
@@ 5.4 Effects of The Exoskeleton on Subjective Feedback
In this study, we also investigated the user experience with the exoskeleton, by quantifying the changes in RPE, RPD, and PW among conditions. The physical demand and effort items in the NASA-TLX and RPE among three conditions presented similar phenomenon as the EMG results. In the HIT-POSE Match condition, they displayed statistical significance from the other two conditions, and the participants perceived the greatest assistance and thus exerted the lowest physical effort. The outcomes are in line with the previous study [12].
In addition, scores of the frustration item in NASA-TLX also exhibited a significant reduction with HIT-POSE conditions, which is in line with the existing study [12]. Evidence may be available in a SUS question (9. I felt very confident using the system.) with a high average score (4.56 ± 0.53). We speculate that mental confidence bridges the frustration from the task, which may be related to a placebo effect. Meanwhile, the SUS score (79.7 ± 5.9) expressed good usability of the exoskeleton according to the empirical evaluation [43].